can someone help me factor these problems? Thanks.
1.4r^3-4
2.(21-7t)+(3r-rt)
3.x^3+64y^3
4.(x^3+7x^2)(-4x-28)

Answer by stanbon(57967)

1.4r^3-4
=4(r^3-1)
=4(r-1)(r^2+r+1)
2.(21-7t)+(3r-rt)
=7(3-t)+r(3-t)
(3-t)(7+r)

3.x^3+64y^3
=(x+4y)(x^2+4xy+16y^2)

4.(x^3+7x^2)+(-4x-28)
=(x^2)(x+7)-4(x+7)
=(x+7))(x^2-4)
=(x+7)(x-2)(x+2)
Cheers,
Stan H.